Please help!!!
What is the percent savings if a stereo is selling for $445.50 and is marked down to $401.95?

Found 2 solutions by RAY100, stanbon:
Answer by RAY100(1637)

savings is 445.50 - 401.95 = 43.55
,
ratio of original = 43.55 / 445.50 = .09775 = 9.8 %
,
or 401.95 / 445.50 = .902,,,, 1 - .902 = .098 = 9.8 %

Answer by stanbon(48535)

What is the percent savings if a stereo is selling for $445.50 and is marked down to $401.95?
---------------------
Let the rate be "r".
Equation:
(1-r)445.5 = 401.95
445.5 - r*445.5 = 401.95
-445.5r = --43.55
e = 0.0978
The discount rate is approximately 9.78%
============================================
Cheers,
Stan H.